OpenStudy (anonymous):

1/t=h+y/x -Solve for x, t, and y?

OpenStudy (anonymous):

there are 3 variables but one equation, i think it can't be solved with the given data.

OpenStudy (across):

Algebraic manipulation

OpenStudy (across):

you have \[\frac{1}{t}=h+\frac{y}{x}\]for t: \[t=\frac{x}{xh+y}\]and so forth

OpenStudy (across):

I believe "solve for" stands for "express in terms of" in this case.

OpenStudy (anonymous):

so this is invalid?

OpenStudy (across):

What is invalid?

OpenStudy (anonymous):

I mean your not actually solving the variable?

OpenStudy (across):

If you're trying to find a number solution, you won't be able to. The Pythagorean theorem \[a^2+b^2=c^2\]cannot be solved as it stands, yet is it "invalid"?

OpenStudy (anonymous):

Nooo I know that but im not looking for a number solution I think what you said by "express in terms" is what im looking for